OpenStudy (anonymous):

Suppose a triangle has sides 3, 4, and 6. Which of the following must be true? A. The triangle in question is a right triangle. B. The triangle in question is not a right triangle. C. The triangle in question may or may not be a right triangle. ***my answer; B is that right? :) @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

how did you get B

OpenStudy (anonymous):

well i'm not sure if i did this right, but i used the pythagorean theorem.. 3^2 + 4^2 = 6^2 9+16 = 36 25 = 36 (which it doesn't!) so i thought that this means its an acute triangle?? not sure if i did that right tho :/

OpenStudy (anonymous):

ah! wait so did i do this correctly?! :O

jimthompson5910 (jim_thompson5910):

you don't need to state whether it's acute or not all you need to do is show that a^2 + b^2 = c^2 is false, which you did when you got to 25 = 36 so that proves this isn't a right triangle

jimthompson5910 (jim_thompson5910):

yes you did, just wanted to make sure you got it and it wasn't a random guess

OpenStudy (anonymous):

ahh! yay!! :) but was i right tho? its acute?! :) and haha yeah no random guesses :P

jimthompson5910 (jim_thompson5910):

idk if it's acute or not, but it's definitely not a right triangle

jimthompson5910 (jim_thompson5910):

if a^2 + b^2 = c^2 is true, then it's a right triangle

OpenStudy (anonymous):

ahh okay!! I see :) Thank youuu!!

OpenStudy (anonymous):

okay :)